OpenStudy (anonymous):

find equations of both lines through the point (2, -3) that are tangent to the parabola y=x^2+x

OpenStudy (dmb1215):

@Mertsj can you help plz

OpenStudy (mertsj):

Take the first derivative of y = x^2+x

OpenStudy (anonymous):

okay then?

OpenStudy (mertsj):

Replace x with 2 to find the slope of the tangent.

OpenStudy (anonymous):

so the slope is 5?

OpenStudy (mertsj):

Oh. I don't think that is right because the given point is not on the parabola.

OpenStudy (anonymous):

i have to find smaller sloper and larger slope

OpenStudy (anonymous):

slope*

OpenStudy (mertsj):

I have to think about this.

OpenStudy (anonymous):

the answer is -x-1 (smaller slope) and 11x-25 (larger slope)

OpenStudy (anonymous):

but i don't understand how

OpenStudy (anonymous):

@Mertsj there?

OpenStudy (mertsj):

ok. pick a point on the parabola. If the x value is x, then y=x^2+x so the point(x,x^2+x) is on the parabola. Find the slope from the given point (2,-3) to (x,x^2+x)

OpenStudy (anonymous):

can i take a point 2 on the parabola?

OpenStudy (mertsj):

That slope is: \[\frac{x^2+x-3}{x-2}\]

OpenStudy (mertsj):

But the slope of the parabola is given by 2x+1 so set the two slopes equal.

OpenStudy (mertsj):

\[\frac{x^2+x-3}{x-2}=2x+1\]

OpenStudy (mertsj):

solve that equation.

OpenStudy (anonymous):

okay

OpenStudy (mertsj):

I made a sign error. In the slope, it should be - (-3)

OpenStudy (mertsj):

\[\frac{x^2+x+3}{x-2}=2x+1\]

OpenStudy (anonymous):

i dont know if im right but i got x=-1 and x=5

OpenStudy (mertsj):

\[x^2+x+3=2x^2-3x-2\] x^2-4x-5=0 (x-5)(x+1)=0 x=5, x=-1

OpenStudy (mertsj):

Do you know what to do now?

OpenStudy (anonymous):

find y?

OpenStudy (mertsj):

yes

OpenStudy (anonymous):

okay give me a sec

OpenStudy (anonymous):

umm ok so i got (-1, 0) and (5, 30)

OpenStudy (mertsj):

Me too.

OpenStudy (anonymous):

now how do i find the smaller and larger slope?

OpenStudy (mertsj):

The points you just found are the points where the tangents from (2,-3) to the parabola intersect the parabola. So now you job is to find the equations of the two lines. One of the lines goes through (5,30) and (2,-3) The other line goes through (-1,0) and (2,-3) The easy way to find the slopes of the two lines is to use the slope function which is the first derivative. y'=m=2x+1 If x = 5, m = 11 If x = -1, m = -1 So now our job is easy. Find the equation of the line whose slope is 11 and contains the point (2,-3) Find the equation of the line whose slope is -1 and contains the point (2,-3)

OpenStudy (mertsj):

\[y+3=11(x-2)\] \[y+3=11x-22\] \[y=11x-25\] And: \[y+3=-1(x-2)\] \[y+3=-x+2\] \[y=-x-1\]

OpenStudy (anonymous):

Thanks a million!!

OpenStudy (mertsj):

yw

OpenStudy (whpalmer4):

The graphical result:

OpenStudy (anonymous):

Thanks

OpenStudy (anonymous):

could you also help me differentiate 5e^x rootx

OpenStudy (whpalmer4):

Chain rule and power rule.

OpenStudy (anonymous):

\[5e ^{x}\sqrt{x}\]

OpenStudy (anonymous):

\[5e ^{x}((1/2)x ^{-1/2})+ \sqrt{x} (5e ^{x})\]

OpenStudy (anonymous):

this doesn't seem right to me

OpenStudy (whpalmer4):

Nevertheless, that is the correct answer.

OpenStudy (anonymous):

\[(1/2)x ^{-1/2}+\sqrt{x}\]

OpenStudy (mertsj):

It is right, however.

OpenStudy (anonymous):

yeah i got it now

OpenStudy (anonymous):

Differentiate (x+9)/(x^3+x-4)

OpenStudy (anonymous):

Please help me out

OpenStudy (anonymous):

\[-(2x^3+27x^2+13)/(x^3+x-4)\]

OpenStudy (anonymous):

i got that

OpenStudy (whpalmer4):

need to square the denominator

OpenStudy (whpalmer4):

appears correct other than that slip

OpenStudy (anonymous):

another question i'm stuck on

OpenStudy (anonymous):

For what values of x does the graph of f have a horizontal tangent (use n as integer variable) Enter answer as a comma separated list. f(x)=x+2sinx

OpenStudy (anonymous):

anybody?

OpenStudy (whpalmer4):

Generally, if medals have already been given out on a given problem, I think it likely that some people won't bother looking at it, presumably feeling that it is unlikely they will get one. So unless there is some connection between the problems, in my opinion you are better off posting a fresh question, rather than continuing on. Some people may also feel that if they are going to answer a bunch of questions, they maybe ought to get more than one medal, too.